OpenStudy (anonymous):

if sin(theta)=-3/5 and theta is in quadrant III, find tan(theta/2)

OpenStudy (campbell_st):

use t = tan (theta/2) and sin(theta) = 2t/(1 + t^2) -3/5 = 2t/(1+t^2) or -3(1+t^2) = 5(2t) simplify 3t^2 +10t + 3= 0 (3t +1)(t + 9) = 0 t = -1/3 and -9

OpenStudy (campbell_st):

-1/3 = tan(theta/2) and -9 = tan(theta/2) ummm something is not quite right

OpenStudy (anonymous):

figured it out...need to use \[\tan \theta/2=(\sin \theta)/(1+\cos \theta)\]